### v2

# OpenRubrics: Towards Scalable Synthetic Rubric Generation for Reward Modeling and LLM Alignment

Abstract
Reward modeling lies at the core of reinforcement learning from human feedback (RLHF), yet most existing reward models rely on scalar or pairwise judgments that fail to capture the multifaceted nature of human preferences. Recent studies have explored rubrics-as-rewards (RaR) that uses structured natural language criteria that capture multiple dimensions of response quality. However, producing rubrics that are both reliable and scalable remains a key challenge. In this work, we introduce OpenRubrics, a diverse, large-scale collection of (prompt, rubric) pairs for training rubric-generation and rubric-based reward models. To elicit discriminative and comprehensive evaluation signals, we introduce Contrastive Rubric Generation (CRG), which derives both hard rules (explicit constraints) and principles (implicit qualities) by contrasting preferred and rejected responses. We further improve reliability by enforcing preference–label consistency via rejection sampling to remove noisy rubrics. Across multiple reward-modeling benchmarks, our rubric-based reward model, Rubric-RM, surpasses strong size-matched baselines by 6.8%. These gains transfer to policy models on instruction-following and biomedical benchmarks. Our results show that rubrics provide scalable alignment signals that narrow the gap between costly human evaluation and automated reward modeling, enabling a new principle-driven paradigm for LLM alignment.
Keywords: Rubrics-as-Rewards, Reward Modeling, LLM Alignment, Synthetic Data
= Date: January 11, 2026
= Model Weights & Checkpoints: https://huggingface.co/OpenRubrics/models
= Datasets: http://huggingface.co/OpenRubrics/datasets
= Contact: liu3351@purdue.edu; ran.xu@emory.edu; hwang28@albany.edu
[…]
1 Introduction
Reward modeling is central to reinforcement learning from human feedback (RLHF) and is widely used to align large language models (LLMs) with human preferences (ouyang2022training, wu2023fine, bhaskar2025language). By assigning a scalar score (ouyang2022training) or preference label (chen2025rm) to each response, reward modeling provides the optimization signal during training and steers the policy LLM toward generating helpful and harmless responses.
While lots of efforts have been paid on RL with verifiable reward (RLVR) (guo2025deepseek, yue2025does), many high-value applications of LLMs, such as long-form question answering, general helpfulness, operate in inherently subjective domains where correctness cannot be sufficiently captured by binary signals. To bridge this gap, rubrics-as-rewards (RaR) (gunjal2025rubrics) have emerged as a new paradigm for reward modeling. Rubrics include structured natural language criteria that decompose quality into interpretable and measurable dimensions, providing a more consistent and transparent evaluation framework than scalar judgments. For policy models, rubrics also enable optimization to be guided by explicit principles.
Despite their great promise, the construction of high-quality rubrics remains an open challenge. Existing benchmarks (arora2025healthbench) curate rubrics with the effort from domain experts, which is costly and difficult to scale. Recent works (huang2025reinforcement, viswanathan2025checklists, gunjal2025rubrics) typically generate rubrics via direct prompting LLMs, but those approaches suffer from limited quality control over rubrics and can be prohibitively expensive when relying on commercial APIs.
In this work, we present OpenRubrics, a large collection of (prompt, rubrics) pairs to facilitate rubric-generation model training. Specifically, we prompt the LLM to generate two complementary types of rubrics: hard rules, which capture explicit and objective constraints specified in the prompt, and principles, which summarize implicit and generalizable qualities of strong responses. This design allows the rubrics to capture both surface-level requirements and deeper dimensions of quality. Although hard rules are typically straightforward to extract, the principles are more subtle and require fine-grained reasoning. To address this, we propose Contrastive Rubric Generation (CRG), which conditions on user queries paired with both chosen and rejected responses. By leveraging negative contrasts, CRG encourages the model to identify discriminative qualities that distinguish stronger answers from weaker ones, yielding more comprehensive and ranking-aware rubric signals. To further ensure reliability and reduce the noise, we apply preference-label consistency through rejection sampling, retaining only rubrics that yield correct preference predictions.
Our contributions are three-fold:
  • We introduce OpenRubrics, a large-scale and diverse collection of rubrics. This dataset enables both rubric generation models and rubric-informed reward modeling at scale.
  • We distinguish between two fundamental types of rubrics and propose a contrastive rubric generation strategy that trains models to produce comprehensive and discriminative rubrics from prompts and responses. Besides, we introduce preference–label consistency that improves the quality and reliability of the rubric.
  • We conduct extensive experiments on eight benchmark datasets, where Rubric-RM consistently outperforms strong baselines by 8.4%. Moreover, when integrated into policy optimization, Rubric-RM consistently yields notable gains on challenging instruction following and medical benchmark. Case studies further verify the benefits of combining hard rules and principles, showing that rubrics help reduce false positives from overly long outputs.
2 Related Works
Reward Modeling.
Standard reward models assign scalar scores to responses by applying a ranking loss under the Bradley–Terry framework (bradley1952rank, liu2025skywork). To enhance reasoning capability, generative reward models (GenRMs) incorporate synthesized Chains of Thought (CoT), enabling more accurate reward estimation (ankner2024critique, yu-etal-2025-self, zhang2025generative, liang-etal-2025-generative). Beyond the pointwise setting, pairwise reward models have been proposed to directly compare multiple responses (xu2025unified, liu2025pairjudge). More recently, reinforcement learning has been leveraged to further optimize reward models, enabling them to reason explicitly over comparisons and thereby achieve stronger alignment performance (chen2025judgelrm, chen2025rm, whitehouse2025j1, guo2025reward). Orthogonal to these efforts, our work focuses on improving reward modeling quality with structured rubrics. By introducing rubric-based evaluation signals, we complement existing approaches with an additional layer of interpretability that yield performance gains.
Rubrics as Rewards.
Recent work has explored rubrics for both evaluation and alignment. Rubrics provide structured assessments of model generations (arora2025healthbench, hashemi2024llm, pathak2025rubric, lin2025wildbench), guide instruction following and domain adaptation (viswanathan2025checklists, gunjal2025rubrics), improve safety via rule-based rewards (mu2024rule), and have been combined with verifiable rewards for reasoning tasks (huang2025reinforcement, zhou2025breaking). Yet most existing approaches rely on prompting frontier LLMs to generate rubrics, which limits scalability and consistency. Our work introduces a more scalable framework for high quality synthetic rubric generation, improving both reward quality and interpretability at a cheaper cost. Concurrently, zhang2025chasing also investigate rubric generation, but focus on iterative refinement to mitigate reward over optimization, whereas we emphasize scalable synthesis and rubric–preference consistency.
3 Preliminaries
Figure 1: Overall Framework for Synthetic Rubric Generation in OpenRubrics.
Rubrics. We define rubrics as a structured set of evaluation criteria tailored to a given prompt. Formally, let $x$ denote an input prompt and $\hat{y}$ a model-generated response. A rubric $\mathcal{R}(x)$ is represented as a collection of $k$ criteria $\mathcal{R}(x)=\{c_{i}\}_{i=1}^{k}$, where each $c_{i}$ denotes a rubric description specifying an aspect of response quality (e.g., factual correctness, reasoning soundness, style).
Rubrics-based Reward Models. Building on prior work in pairwise reward modeling (liu2025pairjudge, chen2025rm, guo2025reward, xu2025unified), we focus on a comparative setting where the goal is to evaluate the relative quality of two candidate responses. Given a prompt $x$ and two samples $(\hat{y}_{1},\hat{y}_{2})$, a pairwise rubric-based reward function is defined as
$\text{reward}_{\text{pair}}(x,\hat{y}_{1},\hat{y}_{2})=r_{\theta}\big(x,\hat{y}_{1},\hat{y}_{2};\{c_{i}\}_{i=1}^{k}\big),$
where reward is the binary preference label, $r_{\theta}$ is the reward model that integrates rubric criteria $\{c_{i}\}$ when producing a preference judgment.
The overall framework for OpenRubrics is in Figure 1. Our overall objective is two-fold: (1) constructing a rubric dataset $\mathcal{D}_{\text{rubric}}$ to train a generation model $g_{\theta}$ that automatically synthesizes rubrics $\mathcal{R}(x)$ given a prompt $x$; and (2) building a reward modeling dataset $\mathcal{D}_{\text{rm}}$ to train a rubric-guided reward model $r_{\phi}$ capable of producing reliable and interpretable pairwise judgments. This two-stage formulation decomposes evaluation into rubric generation and rubric-conditioned reward prediction, bridging human-aligned criteria and automated preference modeling.
4 OpenRubrics
4.1 Data Construction
Data Sources. To generate high-quality rubrics that generalize across tasks and domains, we integrate a range of publicly available preference and instruction-tuning datasets, balancing general-domain data with domain-specific resources. Specifically, our dataset draws from the following sources:
  • UltraFeedback (cui2024ultrafeedback), which aggregates preference annotations from Evol-Instruct (xu2024wizardlm), UltraChat (ultrachat), ShareGPT (sharegpt), and TruthfulQA (lin-etal-2022-truthfulqa).
  • Magpie (xu2025magpie): a large-scale synthetic alignment dataset generated by self-prompting LLMs across diverse domains.
  • Skywork-Preference (liu2024skywork), which integrates data from HelpSteer2 (wang2024helpsteer) and OffsetBias (offsetbias).
  • Synthetic-IF (lambert2025tulu), a collection of human preference judgments tailored for verifiable instruction-following.
  • MegaScience (fan2025megascience), a specialized corpus spanning multiple scientific domains including physics, biology, and chemistry.
  • Medical-o1 (chen2024huatuogpt), a medical SFT dataset curated for diagnostic reasoning tasks.
(a) The data distribution for OpenRubrics (in # instructions).
(b) The distribution for the length of prompts and rubrics, as well as the number of rubrics.
Figure 2: Statistics Overview of OpenRubrics.
Preference Data Construction. To build preference data for rubric generation and judge training (see Sec. 4.3), we reuse existing preference and SFT datasets with tailored processing. For UltraFeedback, we select the highest-scoring response as the chosen and the lowest as the rejected. For MegaScience, and Medical-o1, we generate multiple responses using Qwen-3-8B/14B (yang2025qwen3), Llama-3.1-8B (grattafiori2024llama), and Gemma-3-12B (team2025gemma), selecting one from each model. For Synthetic-IF, responses satisfying all verification functions are labeled as chosen, and others as rejected. For the MegaScience and Medical-o1 datasets, we employ an ensemble of open-source reward models: Athene-RM-8B (Athene2024) and Skywork-Reward-V2-Llama-3.1-8B-40M (liu2025skywork) to rank responses and form best–worst preference pairs.
4.2 Rubrics Synthesis
After collecting a diverse set of preference pairs, our objective is to construct a set of rubrics that serve as anchors to guide reward modeling. To comprehensively represent different types of constraints while preserving discriminative granularity, we categorize rubrics into two types: (1) Hard Rules, which capture explicit requirements stated in the user’s prompt; and (2) Principles, which describe higher-level qualitative aspects such as reasoning soundness, factuality, or stylistic coherence. We then introduce two strategies for generating high-quality rubrics as follows:
Contrastive Rubric Generation. Given a preference dataset $\mathcal{D}=\left\{\left(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-}\right)\right\}_{i=1}^{N},$ where $x_{i}$ is the prompt and $\hat{y}_{i}^{+}$ and $\hat{y}_{i}^{-}$ denote the preferred and displeased responses, respectively. Our objective is to infer rubrics $\mathcal{R}\left(x_{i}\right)$ that capture qualities a good response should satisfy and the criteria by which one response is preferred over another, using $\hat{y}_{i}^{+}$ and $\hat{y}_{i}^{-}$ as guidance. Formally, we prompt an instruction-tuned LLM $h_{\psi}$ as $\mathcal{R}\left(x_{i}\right)\sim h_{\psi}\left(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\ell_{i}\right),$ where $\ell_{i}$ is the preference label. The generator is asked to produce a set of discriminative evaluation criteria $\mathcal{R}\left(x_{i}\right)=\left\{c_{i,1},\ldots,c_{i,k_{i}}\right\}$, where each $c_{i,j}$ describes a specific aspect. This setting encourages the model to discover rubric dimensions that are both task-sensitive and preference-aligned.
Rubric Filtering with Preference-label Consistency. Not all generated rubrics faithfully capture the human preference signal. To ensure reliability, we conduct a consistency-based filtering step via prompting the LLM $h_{\psi}$ again. For each triplet ($x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-}$), we concatenate the full rubric $\mathcal{R}\left(x_{i}\right)$ into the context and ask the model to predict which response better aligns with the rubric $\hat{l}_{i}=h_{\psi}\left(x_{i},\mathcal{R}\left(x_{i}\right),\hat{y}_{i}^{+},\hat{y}_{i}^{-}\right),$ where $\hat{l}_{i}=(\hat{r}_{i},\hat{\ell}_{i})$ denotes the prediction rationale and the predicted preference, respectively. For pairwise judging, we retain only those rubrics that lead to predictions consistent with the original human label $\ell_{i}$:
[…]
For listwise judging settings, where a single prompt is associated with multiple candidate responses, we decompose the task into multiple pairwise comparisons. In this case, we retain a rubric if it yields preference predictions consistent with human labels in at least $50\%$ of the induced pairwise comparisons, allowing robustness to occasional ambiguity while preserving discriminative power.
This yields a collection of high-quality rubrics that are both interpretable and empirically consistent with human preferences. The final rubrics-conditioned preference dataset combines prompts, paired responses, and their verified rubrics as $\mathcal{D}_{\text{rubric }}=\left\{\left(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\mathcal{R}^{*}(x_{i})\right)\right\}_{i=1}^{M}.$
Figure 3: The T-SNE plot for embeddings of prompts.
Rubric Statistics Overview. We analyze the curated rubric set along three axes: (1) domain coverage (instruction following, reasoning, general helpfulness; Figure 2(a)); (2) the balance between hard rules and principles as well as the length of prompts and rubrics (Figure 2(b)); and (3) semantic diversity of prompt topics, visualized via t-SNE on embeddings from Qwen-3-Embedding-0.6B (zhang2025qwen3) (Figure 3). These statistics confirm that the synthesized rubrics provide comprehensive, yet discriminative coverage, forming a foundation for rubric-based reward modeling.
4.3 Reward Model Training and Inference
After collecting the rubrics-based dataset, we proceed to develop a rubric generation model that outputs evaluation rubrics and a reward model Rubric-RM that generates final preference labels.
Rubric Generation. We first fine-tune $g_{\theta}$ to generate rubrics $\mathcal{R}^{*}$ conditioned on the prompt $x$. Given the dataset $\mathcal{D}_{\text{rubric}}=\{(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\mathcal{R}^{*}(x_{i}))\}_{i=1}^{M}$, we optimize the standard cross-entropy objective:
[…]
Reward Model Training. We then train the reward model $r_{\phi}$ to predict preference labels $\hat{l}$ using the generated rubrics. Using $\mathcal{D}_{\text{rm}}=\{(x_{i},\hat{y}_{i}^{+},\hat{y}_{i}^{-},\mathcal{R}^{*}(x_{i}),\hat{l}_{i})\}_{i=1}^{M}$, we optimize $r_{\phi}$ to predict $\hat{l}$ conditioned on the prompt, response pair, and rubric:
[…]
Inference. At inference time, given a pairwise test instance $(x,y^{\text{A}},y^{\text{B}})$, Rubric-RM performs a two-stage process to predict the final preference label: (1) the rubric generator produces (or retrieve a cached) rubric conditioned on the instruction $x$ as $\hat{\mathcal{R}}(x)=g_{\theta}(x)$; (2) the reward model then predicts the verdict over candidate responses $(y^{\text{A}},y^{\text{B}})$ conditioned on the generated rubric from two possible labels $\mathcal{C}=\{\text{A is better},\text{B is better}\}$:
$\hat{l}=\arg\max_{k\in\mathcal{C}}p_{\phi}\big(k\mid x,y^{\text{A}},y^{\text{B}},\hat{\mathcal{R}}(x)\big).$
This ensures that the judgment from Rubric-RM is explicitly grounded in rubric criteria.
5 Experiment
5.1 Datasets and Experiment Settings
Training data. We train both components of Rubric-RM: the rubric generator and the judge, on the curated OpenRubrics as presented in Sec. 4.2. Rubrics are produced with contrastive signals from chosen/rejected responses and filtered by preference–label consistency before use. Unless otherwise noted, we use the science-related slice of OpenRubrics to better match our domain study on HealthBench/medical evaluation.
Backbone and variants. Both the rubric generator and the judge are fine-tuned from Qwen-3-8B (“Rubric-RM-8B”) unless specified. At inference time, Rubric-RM follows the two-stage process as detailed in Sec 4.3. We also report an ensemble variant, voting@5, which aggregates five independently sampled judge trajectories by majority vote.
Baselines. We compare Rubric-RM against strong, same-scale white-box reward (judge) models: JudgeLRM-7B (chen2025judgelrm), RRM-7B (guo2025reward), and RM-R1-7B (chen2025rm). We also report larger RM-R1-14B (chen2025rm) and API judges for reference when available. To isolate the benefit of rubric-aware fine-tuning, we include naive Qwen-3-8B (Rubric+Judge) that directly prompts the base model to produce rubrics and then make judgments.
Evaluation benchmarks and metrics. We evaluate Rubric-RM as a pairwise reward model on popular reward-modeling benchmarks: RewardBench (Chat, Chat-Hard) (rewardbench), RM-Bench (liu2025rmbench), PPE-IFEval (ppe), FollowBench (followbench), InfoBench (infobench), IFBench (ifbench), and RewardBench2 (Precise-IF, Focus) (malik2025rewardbench2). While FollowBench and InfoBench were originally designed to assess instruction-following capabilities of LLMs, we adapt them to pairwise evaluation settings by sampling two responses from the same model (Qwen-3-8B/14B), where one response adheres to all specified constraints and the other violates some of them. For the domain study we additionally report HealthBench/medical results. We follow each benchmark’s official splits and scoring rules, reporting accuracy, win-rate or other specific scores.
Due to space limits, additioanl implementation details are deferred to Appendix B.
RewardBench IF Evaluation Benchmarks RM-Bench RewardBench2 HelpSteer3 Avg.
Chat Chat Hard FollowBench PPE-IFEval InfoBench IFBench Chat Precise IF Focus
Black-box LLMs (For reference only)
Claude-3.5-Sonnet 96.4 74.0 58.0 62.5 38.8 87.0 -
API (Rubric+Judge) 79.6 79.2 83.2 61.0 82.2 66.2 67.9 42.5 79.6 71.4 71.3
API (direct Judge) 89.6 71.2 81.7 59.2 72.9 60.4 67.2 13.2 63.4 70.3 64.9
Larger White-box LLMs (For reference only)
RM-R1-14B (Qwen-2.5-Inst) 73.5 79.8 84.0 59.0 85.5 60.8 73.2 23.8 84.6 74.8 69.9
RM-R1-14B (DeepSeek-Dist) 90.3 78.9 89.9 61.2 82.4 59.0 71.4 30.6 79.0 74.6 71.7
White-box Judge/Reward LLMs
JudgeLRM-7B 92.1 56.1 79.8 46.0 62.7 47.5 55.4 9.4 29.1 60.2 53.8
RRM-7B 77.7 69.5 65.5 51.0 68.2 53.2 59.9 10.0 60.4 62.4 57.8
RM-R1-7B (Qwen-2.5-Inst) 83.0 70.0 56.3 55.2 71.3 55.2 64.2 20.6 76.2 65.2 61.7
RM-R1-7B (DeepSeek-Dist) 85.3 67.3 69.7 51.0 70.3 56.5 62.2 13.8 55.4 62.6 59.4
Qwen-3-8B (Rubric+Judge) 74.2 64.2 71.3 57.0 74.3 59.5 63.9 8.1 44.0 60.8 57.7
Rubric-RM (Our proposed model)
Rubric-RM-4B 87.2 68.9 78.0 64.8 79.3 63.2 62.6 35.6 79.6 64.7 68.4
Rubric-RM-4B-voting@5 88.7 70.2 80.7 66.2 83.0 64.9 63.8 38.1 82.6 65.0 70.3
Rubric-RM-8B 88.2 74.1 76.1 67.0 80.8 65.4 65.7 34.4 82.2 67.0 70.1
Rubric-RM-8B-voting@5 89.9 75.4 81.5 70.8 83.8 67.1 67.0 40.0 86.5 67.5 73.0
Table 1: Comparison of different judge and reward models across multiple benchmarks. RewardBench2 reports results on Precise IF, and Focus dimensions. Rubric API uses GPT-4.1-Mini, and Judge API uses Gemini-2.5-Flash-Lite. Best results are highlighted in bold.
5.2 Performance of Rubric-RM
We first validate the performance of Rubric-RM for reward modeling. For a more systematic evaluation, we test both 4B and 8B variants of Rubric-RM, which use Qwen3-4B/8B as backbones. Table 1 reports the results of Rubric-RM.
Outperforming Comparable Reward Models. Both Rubric-RM-4B and 8B surpass all existing 7B-scale white-box baselines (e.g., JudgeLRM, RRM, RM-R1). Notably, Rubric-RM-4B (68.4) outperforms the strongest 7B competitors (max 61.7), with Rubric-RM-8B further improving to 70.1. This confirms that rubric-aware training yields more reliable signals than generic preference learning, even with reduced parameter counts.
Majority Voting Further Enhances Performance. We also evaluate Rubric-RM-voting@5, which aggregates predictions via majority voting across five independent judge trajectories. Rubric-RM-4B-voting@5 reaches 70.3, and Rubric-RM-8B-voting@5 achieves the best overall average of 73.0, surpassing much larger models such as RM-R1-14B (71.7) and the Rubric+Judge API (71.3). These results highlight the stability benefits of rubric-based ensembles.
Effectiveness of Rubric-Aware Fine-Tuning. Directly using Qwen-3-8B to generate rubrics and judge yields only 57.7. In contrast, Rubric-RM achieves 70.1. This +12.4 gain validates our core contribution: high-quality rubrics derived via contrastive generation and consistency filtering are essential for effective reward modeling.
Strength on IF Benchmarks. Rubric-RM excels on benchmarks requiring fine-grained instruction adherence. On FollowBench and InfoBench, we achieve 81.5 and 83.8 respectively, substantially outperforming baselines like JudgeLRM and RRM. This demonstrates that rubrics capture nuanced constraints better than scalar reward models.
In the remaining experiments, we use Rubric-RM-8B as our reward model unless specified.
Model IFEval (Prompt) IFEval (Inst.) IFEval InfoBench
Loose Strict Loose Strict AVG AVG
GPT-4 (0314) 79.3 76.9 85.4 83.6 81.3 87.3
AutoIF (dong2024self) 56.9 47.1 67.0 57.6 57.2 80.6
UltraIF (an2025ultraif) 75.4 71.3 83.0 79.4 77.3 80.7
Qwen2.5-7B-Instruct 75.0 72.5 81.8 79.9 77.3 78.1 (76.0)
+ SFT (Distilled) 66.8 64.1 75.3 72.8 69.8 72.5
+ DPO (via Skywork) 75.7 68.0 83.2 78.5 76.0 82.0
+ DPO (via ArmoRM) 73.8 70.2 81.7 78.3 76.0 83.5
+ DPO (via Ultrafbk.) 71.5 69.1 79.9 77.7 74.6 80.0
+ DPO (via AI Judge) 73.0 68.9 80.9 77.8 75.2 76.1
+ DPO (via RLCF) 77.3 72.6 84.1 80.3 78.6 84.1 (81.5)
+ DPO (via Rubric-RM) 78.2 73.9 84.5 81.2 79.5 83.0
Table 2: Comparison of trained policy models with different reward models on a format-based instruction-following benchmark (IFEval) and an open-ended benchmark (InfoBench). Baseline results are from viswanathan2025checklists. Results with underlines are reproduced by us using official checkpoints and evaluation scripts. Best scores are in bold.
Model Arena-Hard AlpacaEval AVG
Vanilla SC Vanilla LC
GPT-4 (0314) 50.0 50.0 22.1 35.3 39.4
UltraIF (an2025ultraif) 31.4
Qwen2.5-7B-Instruct 51.3 42.8 33.5 36.2 41.0
+ SFT (Distilled) 32.6 29.2 36.1 33.3 32.8
+ DPO (via Skywork) 55.1 50.3 44.8 41.5 47.9
+ DPO (via ArmoRM) 50.8 46.4 37.6 38.1 43.2
+ DPO (via Ultrafbk.) 52.8 47.9 33.7 38.7 43.3
+ DPO (via AI Judge) 51.0 44.4 28.8 33.4 39.4
+ DPO (via RLCF) 54.6 48.4 36.2 37.1 44.1
+ DPO (via Rubric-RM) 52.9 53.1 47.0 41.3 48.6
Table 3: Comparison of different alignment strategies applied to Qwen2.5-7B-Instruct on Arena-Hard and AlpacaEval. Baseline results are from viswanathan2025checklists. SC and LC stands for Style-controlled and Length-controlled. Best results are in bold.
Figure 4: Comparison of policy models on IFBench. Results of baselines except RLCF are from bhaskar2025language. We evaluate RLCF with its official checkpoint.
Method Creative Planning Math Info seeking Coding WB Score
Claude-3.5-Sonnet 55.6 55.6 50.2 55.5 56.5 54.7
GPT-4-turbo 58.7 56.2 51.0 57.2 55.1 55.2
GPT-4o-mini 60.1 58.2 54.0 57.4 57.2 57.1
Qwen2.5-7B-Instruct 50.1 51.8 47.1 50.7 45.0 48.7
+DRIFT 52.5 53.2 50.6 52.4 50.3 51.7
+DPO (via RLCF) 51.4 52.7 49.0 51.3 48.8 50.5
+DPO (via RLMT (PPO)) 52.1 52.6 45.2 51.4 48.3 49.7
+DPO (via Rubric-RM) 54.8 55.5 51.5 54.1 52.9 53.6
Table 4: Comparison of different alignment strategies applied to Qwen2.5-7B-Instruct on WildBench. Results are reported for task-specific scores and task macro WB score. Baseline results with "*" are from wang2025drift. We evaluate RLCF and RLMT with their official checkpoints. Best results are in bold.
5.3 Policy Models with Rubric-RM
5.3.1 Instruction-Following Evaluation
We further evaluate the effectiveness of using Rubric-RM as a reward model for policy optimization on instruction-following tasks, including IFEval, InfoBench, and IFBench. The results are shown in Table 2 and Figure 4.
Improved Performance on IFEval and InfoBench. Using Rubric-RM yields the best overall performance among open-source reward-model baselines: the resulting policy model reaches 79.5 on IFEval (vs. 76.0 with Skywork/ArmoRM) and 83.0 on InfoBench, approaching much larger commercial systems. These results indicate that rubric-based rewards provide more reliable optimization signals for constrained instruction following.
Clear Gains on Complex Instruction Following (IFBench). Figure 4 shows that the policy model optimized with Rubric-RM achieves 30.3 on IFBench, substantially higher than RLCF (28.2) and RLMT-based methods (22.4–22.8). Compared with both supervised fine-tuning variants and reinforcement learning baselines, Rubric-RM provides stronger inductive biases, enabling policies to better capture fine-grained instruction adherence.
Overall, these results demonstrate that Rubric-RM, when used as a reward model, provides a substantially stronger training signal that improves the instruction-following capability of learned policies.
5.3.2 General Alignment Evaluation
We evaluate policy model trained with Rubric-RM on alignment benchmarks Arena-Hard, AlpacaEval, and WildBench (Tables 3 and 4).
With DPO optimization, Rubric-RM achieves the best overall average score (48.6) among all open-source reward models. On Arena-Hard (style-controlled), it obtains 53.1, outperforming Skywork (50.3), Ultrafeedback (47.9), and RLCF (48.4). On AlpacaEval, it reaches 47.0, surpassing Skywork (44.8) and ArmoRM (37.6). On WildBench, aligning Qwen2.5-7B-Instruct with Rubric-RM yields the best macro WB score of 53.6, improving over the base model (48.7) by +4.9 and outperforming DPO with RLCF (50.5), RLMT (49.7), and DRN (51.7), with consistent gains across all task categories. These results show that rubric-based signals provide reliable gains across both vanilla and controlled settings.
5.4 Rubric-RM for BioMedical Domain
In this section we further study the effectiveness of Rubric-RM in more specialized biomedical domain, following arora2025healthbench. The Rubric and Judge are Qwen-3-8B backbones fine-tuned with OpenRubrics data from science-related domains.
Performance on HealthBench. Rubric-RM outperforms comparable-size reasoning reward models on HealthBench, achieving 68.3, exceeding RRM-7B (63.3) and RM-R1-7B variants (55.4/66.9), and approaching RM-R1-14B (69.9). Majority voting further boosts performance: Rubric-RM-voting@5 reaches 72.9, narrowing the gap to 14B reasoning models and API-based judges. A key finding is the importance of domain- specific rubric SFT. Directly prompting Qwen-3-8B with a Rubric+Judge pipeline yields only 51.8, whereas Rubric-RM improves performance by +16.5. This highlights our core contribution: contrastive, domain-tuned rubric training produces higher-precision, science-aware evaluation signals than on-the-fly rubric generation.
(a) Judges and Reward Models.
(b) Trained Policy Models.
Figure 5: Comparison of different judges, reward models, and trained policy models on HealthBench.
Preference Optimization with Rubric-RM on HealthBench. We use Rubric-RM as the preference judge for DPO on HealthBench. Starting from Qwen-2.5-7B-Instruct (21.6), DPO with ArmoRM and RM-R1-7B improves performance to 22.5 and 22.7, respectively. In contrast, DPO with Rubric-RM achieves the best result at 23.8, yielding a consistent +1.1–1.3 gain over strong 7B reasoning rewards. These results demonstrate that rubric-aware, domain-specialized reward models translate directly into stronger biomedical policy learning, outperforming generative reasoning rewards.
Case Study on RewardBench Chat Hard
Prompt Describe a vivid and unique character, using strong imagery and creative language. Please answer in fewer than two paragraphs.
Resp A (snippet) “She was a woman of great power and influence …with a deep voice and piercing eyes …” (single paragraph, on-task)
Resp B (snippet) “…a fierce and determined young woman …[paragraph break] …Which of the following is an example of a character …[paragraph break] A character …Lily …”
Label Resp A.
RRM-7B “…both responses are within that limit …Assistant 2’s response is more detailed …(Choose B)
JudgeLRM “…Assistant 2 provided a more vivid description …(Choose B)
Rubric-RM
Rubric “1. Describe a vivid and unique character. [Hard Rule] 2. Use strong imagery and creative language. [Hard Rule] 3. Be fewer than two paragraphs. [Hard Rule] 4. Employ sensory-rich language. [Principle] 5. Use original expressions. [Principle] 6. Maintain a consistent tone. [Principle] 7. Remain cohesive and focused. [Principle] 8. Avoid unnecessary elaboration. [Principle]”
Judge “The paragraph limit is the most verifiable constraint. Response B fails the paragraph rule, while Response A satisfies all hard rules and most principles …(Choose A)
Table 5: Case study with error highlighting. Baselines favor the longer, imagery-rich response and miss the explicit paragraph constraint, while Rubric-RM enforces hard rules before evaluating principles.
5.5 Efficiency Comparison
Compute Time (sec.)
JudgeLRM-7B 25.71
RRM-7B 203.4
RM-R1-7B (Qwen-2.5-Inst) 260.37
RM-R1-7B (DeepSeek-Dist) 170.76
RM-R1-14B (Qwen-2.5-Inst) 322.79
RM-R1-14B (DeepSeek-Dist) 382.02
Rubric-RM-8B 130.77
Table 6: Computing speed on 100 samples (vLLM).
Table 6 reports wall-clock time on 100 randomly sampled prompts from RewardBench2. Despite using two Qwen-3-8B components (rubric generator + judge), Rubric-RM runs in 130.77s, which is no slower than existing reasoning reward models, including RRM-7B (203.4s) and RM-R1-7B/14B (170.8–382.0s). It is substantially faster than 14B R1 variants and competitive with strong 7B reasoning baselines. Efficiency stems from our architecture: rather than long Chain-of-Thought decoding, we use short rubric generation followed by a lightweight judge. Moreover, rubrics are amortizable and can be cached for reuse across many examples, removing rubric generation cost during large-scale scoring. Although JudgeLRM achieves lower latency, it lacks the explicit, interpretable signals necessary for downstream policy optimization.
5.6 Case Studies
We conclude this section with a case study that illustrates how Rubric-RM handles challenging inputs and improves reward modeling. Table 5 shows the case study on RewardBench, where both responses are vivid, but the instruction requires fewer than two paragraphs. The baselines miss this hard constraint and incorrectly prefer the longer answer, exhibiting a verbosity bias and an instruction violation. In contrast, Rubric-RM first applies a simple structural check to reject the non-compliant candidate, then compares higher-level qualities (e.g., imagery, originality, focus), and selects the correct response. This example shows how long CoT can still overlook explicit constraints, while rubric-based decomposition makes such failures clear.
6 Conclusion
We present OpenRubrics, a large-scale dataset and framework for scalable and high-quality rubric generation. By decomposing evaluation into hard rules and principles through Contrastive Rubric Generation and applying preference–label consistency filtering, we construct interpretable and discriminative rubric signals that better align with human judgment. Our rubric-based reward model, Rubric-RM, delivers an average 8.4% improvement across benchmarks and further boosts policy performance on diverse benchmarks with offline reinforcement learning. These results position rubrics-as-rewards as a practical foundation for transparent and generalizable LLM alignment.
Limitations
While OpenRubrics demonstrates strong empirical gains, several limitations remain. First, our rubric generation relies on contrastive signals from preference data and inherited model judgments. Although preference-label consistency filtering reduces noise, the resulting rubrics may still reflect biases present in the underlying models and datasets, particularly for subjective or culturally nuanced criteria. Second, our framework focuses on pairwise comparative evaluation; extending rubric-based rewards to absolute scoring or multi-response ranking scenarios remains an open challenge. Finally, we primarily evaluate rubric-based rewards in offline preference optimization. How such structured rewards interact with fully online RLHF pipelines, exploration dynamics, and long-horizon policy learning remains an important direction for future work.
Appendix A Additional Rubric Statistics
We present additional rubric statistics in the number of preference pairs in Figure 6.
Appendix B Additional Implementation Details
B.1 Decoding and efficiency protocol
All models are run under matched decoding budgets (temperature, max tokens, and stop conditions per benchmark recommendations). We use unified execution stack vLLM (kwon2023efficient) for throughput-fair comparisons. For efficiency (Table 6), we measure wall-clock time to score a fixed set of prompts; note that rubrics from stage (i) are cacheable and can be reused across examples, amortizing the cost in large-scale judging and preference optimization.
B.2 Hyper-parameters
Table 7 details the hyper-parameters used in Rubric-RM and policy model training, which were conducted in LLaMA-Factory (zheng2024llamafactory). Moreover, Table 8 presents sampling parameters used in OpenRubrics curation and Rubric-RM inference. For baseline methods, we adopted the sampling parameters from their official implementations and papers.
Parameter Value
Rubric-RM SFT
Rubric-Generator Epochs 1
Cutoff Length 3072
Batch Size 128
Optimizer AdamW
Learning Rate $8\times 10^{-6}$
LR Schedule Cosine
Warmup 0.05
Judge-Generator Epochs 2
Cutoff Length 6144
Batch Size 64
Optimizer AdamW
Learning Rate $5\times 10^{-6}$
LR Schedule /
Warmup /
Policy Model DPO
Policy Model Epochs 1
Cutoff Length 2048
Batch Size 64
Optimizer AdamW
Learning Rate $3\times 10^{-7}$
LR Schedule /
Warmup /
SFT mixing weight 0.1
$\beta$ 0.1
Table 7: Hyper-parameters used in Rubric-RM and policy model training.
Parameter Value
OpenRubrics Curation
Rubric-Generator Model GPT-4.1-Mini
Maximum Tokens 768
Temperature 0.0
Top-$P$ /
Top-$K$ /
Judge-Generator Model Gemini-2.5-Flash-Lite
Maximum Tokens 2048
Temperature 0.0
Top-$P$ /
Top-$K$ /
Rubric-RM Inference
Rubric-Generator Base-Model Qwen-3-4B/8B (Default)
Maximum Tokens 1024
Temperature 0.0
Top-$P$ /
Top-$K$ /
Enable-thinking False
Judge-Generator Model Qwen-3-4B/8B (Default)
Maximum Tokens 4096
Temperature 0.7
Top-$P$ 1.0
Top-$K$ -1 (All)
Enable-thinking False
Table 8: Sampling parameters used in OpenRubrics curation and Rubric-RM inference.
Case Study on FollowBench
Prompt Would you consider direct air carbon capture as a significant financial venture? In addition to sharing your perspective, incorporate a quote from a recent news article or study to substantiate your viewpoint, and concisely summarize it. Also, ensure to mention the publication date of the referenced source. Furthermore, briefly discuss the potential economic implications based on your source.
Resp A (snippet) “Direct air carbon capture (DAC) is a promising technology …According to a recent study published in Nature Energy in March 2023 …In conclusion, while there are still challenges to overcome, the recent study suggests that DAC has the potential to become a financially viable option.”
Resp B (snippet) “‘…Would you consider direct air carbon capture as a significant financial venture? Yes, I absolutely believe …a recent report from BloombergNEF, published on May 16, 2024, titled …BNEF’s projection highlights several key economic implications: 1. …”
Ground Truth Resp B.
RM-R1-7B (DS) “ …looking at Chatbot A’s response. It starts by explaining DAC briefly and then moves into the study from Nature Energy in 2023 …Chatbot B’ …The quote is from BloombergNEF, but the publication date isn’t provided(Choose A)
RM-R1-7B (Qwen) “ …Chatbot A follows all instructions by providing a quote from a recent study …Chatbot B appears to have misunderstood the prompt …doesn’t provide a proper citation with the requested publication date(Choose A)
Rubric-RM
Rubric “1. The response must explicitly state whether direct air carbon capture is considered a significant financial venture. [Hard Rule] 2. The response must include a direct quotation from a recent news article or study. [Hard Rule] 3. The response must provide a concise summary of the quoted source. [Hard Rule] 4. The response must mention the publication date of the referenced source. [Hard Rule] 5. The response must briefly discuss the potential economic implications based on the source. [Hard Rule] 6. The response should integrate the cited evidence seamlessly into the argument to support its claims. [Principle] 7. The response should maintain a clear and logical structure, guiding the reader through the argument. [Principle] 8. The response should use precise and concise language, avoiding unnecessary verbosity. [Principle] 9. The response should ensure that all supporting details are directly relevant to the central argument. [Principle] 10. The response should present a balanced analysis by acknowledging both potential benefits and challenges. [Principle]”
Judge “The most objective rule is the requirement to explicitly state whether direct air carbon capture is considered a significant financial venture …Response A fails …not explicitly stating whether DAC is(Choose B)
Table 9: FollowBench case study with error highlighting. Baselines pick A due to a hallucinated “missing date” and missed stance requirement; RUBRIC-RM correctly picks B.
B.3 Additional Case Study
Another example is shown in Table 9. This example is more challenging: both answers are longer and the quality gap is subtle. Baselines nevertheless produce factual mistakes about the evidence, e.g., asserting that the better response lacks a date/citation, despite it correctly providing a BloombergNEF quote with a May 16, 2024 publication date and concrete figures (\$387B cumulative investment). Our rubric-aware judge identifies recency and verifiability as hard requirements (quote, date, concise summary, and economic implications), and favors the response that meets them. This demonstrates Rubric-RM’s robustness to citation hallucinations and over-weighting of “academic-looking” prose that misled generative reasoning RMs.
B.4 Prompts
We present the prompts we used in this subsection. For baseline methods, we adopted the prompts from their official implementations and papers.
Prompt for Listwise Contrastive Rubric Generation (OpenRubrics Curation) Prompt for Pairwise Contrastive Rubric Generation (OpenRubrics Curation)
Prompt for General Domain Judge Generation (OpenRubrics Curation)
Prompt for Medical Domain Judge Generation (OpenRubrics Curation)
Prompt for Rubric Generation (Rubric-RM)
Prompt for General Domain Judge Generation (Rubric-RM)
Prompt for Medical Domain Judge Generation (Rubric-RM)